# Beyond Release: Access Considerations for Generative AI Systems

### v1

@@ 4.1.4 Comparing models
Comparing hosting and inference costs for open-weight models depends highly on utilization, task, latency, and hardware used, and varies heavily. Given numbers are broad approximations, and change with hardware, usage, and other factors. Precision affects cost, with many popular models trained with 32 or 16 bits of precision. Memory needed for inference tends to be around 20% higher than model memory [26]. While hardware memory varies, we use NVIDIA H100 GPUs, which have 80 GB of VRAM, and NVIDIA H200 GPUs, which have 141 GB of HBM3e memory. These GPUs cost approximately \$25,000 and \$32,000 each, respectively. Costs for hardware, renting GPUs, energy, and production models fluctuate heavily over time. See Appendix Resourcing Calculations for more information on determining resourcing for Llama 3.1 405B Instruct..
@@ 8 Conclusion
Aspects of access beyond system release provide more clarity to release benefits, risks, and tradeoffs. Examining system accessibility and possible scale of access can more accurately inform release and policy decisions and research. We show through granular analysis and model comparisons that similar considerations exist across the release spectrum and for both open- and closed-weight models. We also show barriers without usable infrastructure, and high risk via accessible interfaces and fine-tuning. As accessibility positively impacts scale, reaching more users broadens usage but also reaches more malicious actors and affects ability to intervene. The resourcing, technical usability, and utility subsections of access exemplify how to more meaningfully weigh usefulness and threats.
@@
Thank you to Hugging Face for supporting this research.

### v2

@@ 4 Breaking Down Access
While these three categories necessarily overlap, this framing better structures the core means of using a system. Resourcing applies most to hardware and monetary costs of accessing a system. Technical usability refers most to software features that enable usage. Utility covers qualities and capabilities such as language and modality of system components, which also overlaps with but is distinct from task performance. Some variables, such as ability to adapt or fine-tune, can apply to multiple categories (Resourcing and Technical Usability), but are sorted into the most applicable category. The three categories ensure that access is considered from three vantage points.
Risks and benefits affect stakeholders throughout the deployment timeline, such as developers and deployers, third-party deployers who are not the developers, researchers, users and customers, policymakers and regulators, and members of society who are affected by AI outputs and distribution. Risks and benefits noted are non-exhasutive, and do repeat throughout variables and access categories. This repetition stresses how many access variables can affect accessibility and have similar outcomes. Listed risks and benefits are tied to the listed variable, not necessarily deployer decisions. Stakeholders affected are be different by variable and scenario; for example, some models’ inference cost may disproportionately affect researchers compared to third-party deployers. How tradeoffs are weighed will depend on deployers determining release and policymakers’ priorities.
@@ 4.1.4 Comparing models
Comparing hosting and inference costs for open-weight models depends highly on utilization, task, latency, and hardware used, and varies heavily. Given numbers are broad approximations, and change with hardware, usage, and other factors. Precision affects cost, with many popular models trained with 32 or 16 bits of precision. Memory needed for inference tends to be around 20% higher than model memory [26]. While hardware memory varies, we use NVIDIA H100 GPUs, which have 80 GB of VRAM, and NVIDIA H200 GPUs, which have 141 GB of HBM3e memory. These GPUs cost approximately \$25,000 and \$32,000 each, respectively. Costs for hardware, renting GPUs, energy, and production models fluctuate heavily over time. See Appendix Resourcing Calculations for more information on determining resourcing for Llama 3.1 405B Instruct.
@@ 8 Conclusion
Aspects of access beyond system release provide more clarity to release benefits, risks, and tradeoffs. Examining system accessibility and possible scale of access can more accurately inform release and policy decisions and research on release. We show through granular analysis and model comparisons that similar considerations exist across the release spectrum and for both open- and closed-weight models. We also show barriers without usable resources and infrastructure, and high risk via accessible interfaces and fine-tuning. As accessibility positively impacts scale, reaching more users broadens usage but also reaches more malicious actors and affects ability to intervene. The resourcing, technical usability, and utility subsections of access better exemplify how to more meaningfully weigh system usefulness and threats.